OpenStudy (anonymous):

Find the relationship between x and y so that (x,y) is equidistant (equal distance) from the two points (4,-1) and (-2,3)

OpenStudy (anonymous):

what is this asking us to do exactly?

OpenStudy (amistre64):

find the midpoint?

OpenStudy (anonymous):

idk...

OpenStudy (amistre64):

2 points define a line; so the best option is midpoint id say

OpenStudy (amistre64):

or; the line that runs thru them.. thats it

OpenStudy (anonymous):

perpendicular distance?

OpenStudy (amistre64):

find slope of the 2 points; find the midpoint; and determine the line equation that is perp

OpenStudy (amistre64):

that runs thru the midpoint

OpenStudy (anonymous):

what formula would i use for that

OpenStudy (amistre64):

slope formula to determine slope; midpoint formula to determine midpoint properties of perp lines to determine slope of perp line and slope point form to determine equation of line thru midpoint

OpenStudy (amistre64):

(4,-1) -(-2,3) ------ (6,-4); slope = y/x = -4/6 = -2/3 perp slope = 3/2

OpenStudy (anonymous):

ok thankyou

OpenStudy (amistre64):

midpoint is add em and divide by 2 (4,-1) (-2,3) ------- (2,2) ; /2 = (1,1)

OpenStudy (amistre64):

y-1 = (3/2)(x-1)

OpenStudy (amistre64):

got no idea what form of the line equation they are hunting for tho

OpenStudy (amistre64):

slope intercept form would be: y = (3/2)x -(3(1)/2) +1 y = (3/2)x -1.5 +1 y = (3/2)x - (1/2`)

OpenStudy (amistre64):

without the tilde of course :) fat fingers and tiny keyboard dont mix

OpenStudy (anonymous):

:S huh? lol ok thanks so much amistre